OpenStudy (anonymous):

Solve 3/7 + 5/9. Then use estimation to check for reasonableness. Show your work.

OpenStudy (anonymous):

Find L.C.M of 7 and 9

OpenStudy (anonymous):

did u?

OpenStudy (anonymous):

@Mommasboy u there?

OpenStudy (anonymous):

\[\frac{a}{b}+ \frac{c}{d}=\frac{ad+bc}{bc}\]